Find two consecutive even whole numbers whose product is 224
Answer by mananth(16949)

let he first number be x
the second will be x+2
their product = 224
x(x+2)=224
x^2+2x=224
x^2+2x-224=0
x^2+16x-14x-224=0
x(x+16)-14(x+16)=0
(x-14)(x+16)=0
x=14 OR -16
the numbers can be 14, 16
or
-16, -14